OpenStudy (sphott51):

Planning ahead for Katie's freshman year of college, her parents invested $8,000 into an account at 6% to be compounded annually. After how many years will the account be worth $16,098 ? I used the formula A(t) = A(1 +(r/))^nt

OpenStudy (sphott51):

I think that's the right formula

OpenStudy (alexrobin13):

I'm not sure if you just mistyped the formula but it is A(t) = A(1 +(r/n))^nt

OpenStudy (alexrobin13):

Plug in the values and let me know if you get the answer youre looking for

OpenStudy (sphott51):

I plugged in 16,098 = 8,000(1 + (.06/1)) ^1t

OpenStudy (erak):

Okay, so what do you think would be your first steps for isolating t?

OpenStudy (alexrobin13):

divide 8000 over and go from there

OpenStudy (sphott51):

I divided it and I got 2.01

OpenStudy (erak):

Okay the way I would do it would be with logarithms, like so: \[\frac{ \log2.01 }{\log 1.06 }\] You should get approximately 12 years.

OpenStudy (sphott51):

It isn't logs?

OpenStudy (sphott51):

Oh nevermind haha

OpenStudy (erak):

The function is called a logarithm. It's "log" on a calculator.

OpenStudy (sphott51):

I got 11.98

OpenStudy (erak):

Yeah, so if you didn't round and used 2.01225 instead, you would have gotten basically 12 years. You can double check that with the original equation.

OpenStudy (sphott51):

I could round it with the .9 of 11.98

OpenStudy (erak):

Yeah, just list your answer as "approximately 12 years."